Assume that the weight loss for the first month of a diet program varies between 6 pounds and 12 pounds, and is spread evenly over the range of possibilities, so that there is a uniform distribution. Find the probability of the given range of pounds lost.
More than 10 pounds?

Answer by stanbon(75887)

Assume that the weight loss for the first month of a diet program varies between 6 pounds and 12 pounds, and is spread evenly over the range of possibilities, so that there is a uniform distribution. Find the probability of the given range of pounds lost.
More than 10 pounds?
--------
Draw the picture.
Place the numbers 6 to 12 in order on a horizontal axis.
That stretch is 6 wide.
So the probability over each 1 lb. interval is 1/6
----
The interval from 10 to 12 is a 2 lb interval
The probability over this interval is 2(1/6) = 2/6 = 1/3
==========================================================
Cheers,
Stan H.